OpenStudy (anonymous):

Would someone please help me break down the steps of proving the following:? 1+cos^2(-x)/1-csc^2(-x) = -sin^2x-tan^2x

OpenStudy (anonymous):

Is this \[\frac{1+cos^2{x}}{1-\csc^2{x}}=\sin^2{x}-\tan^2{x}\] ? Change everything into sin and cos and then start mass reducing if you can

hero (hero):

He has negative x's in the parentheses

OpenStudy (anonymous):

I've tried but I get stuck... And yes those are negative x's

OpenStudy (anonymous):

\[\frac{1+cos^2{(-x)}}{1-\csc^2{(-x)}}\] \[\frac{1+cos^2{(-x)}}{1-\frac{1}{\sin^2{(-x)}}}\] \[\frac{1+cos^2{(-x)}}{1-\frac{1}{\sin^2{(-x)}}}\cdot \frac{\sin^2{(-x)}}{\sin^2{(-x)}}\] \[\frac{\sin^2{(-x)}+\sin^2{(-x)}cos^2{(-x)}}{\sin^2{(-x)}-1}\] \[\frac{\sin^2{(-x)}+\sin^2{(-x)}cos^2{(-x)}}{-\cos^2{(-x)}}\] \[-\frac{\sin^2{(-x)}}{\cos^2{(-x)}}-\sin^2{(-x)}\] \[-\tan^2{(-x)}-\sin^2{(-x)}\] Note that sin(-x)=-sin(x) (-sin(x))^2 = sin(x) so \[-\tan^2{(x)}-\sin^2{(x)}\]

OpenStudy (anonymous):

oh and \[1=\sin^2{x} + \cos^2{x}\] so \[\sin^2{x} - 1=-\cos^2{x}\]

OpenStudy (anonymous):

and \[\tan{(-x)} = -\tan{(x)}\] so \[\tan^2{(-x)} = \tan^2{(x})\]

OpenStudy (anonymous):

Oh and above that should be \[\sin^2{(-x)}=\sin^2{(x)}\] I omitted the square term

OpenStudy (anonymous):

Thank you